Write the equation of the line containing the points (-4, 1) and (5, 2).
Answer by mananth(12270)

equation of line passing through two points (x1,y1) & (x2,y2)
=y-y1/x-x1 = y2-y1/x2-x1
The points are(-4,1) (5,2)
..
y-1 / x+4 = 2-1 / 5+4
y-1 /x+4 = 1/9
9(y-1)=1(x+4)
9y-9=x+4
9y-x=13